# InternVideo2: Scaling Video Foundation Models for Multimodal Video Understanding

### v1

@@ 3.3 Stage3: Predicting Next Token with Video-Centric Inputs
To further enrich the semantics embedded in InternVideo2 and improve its user-friendly in human communications, we tune it by incorporating it into a VideoLLM with a LLM and a video BLIP Li et al. (2022b, c). We employ the progressive learning scheme in Li et al. (2023b) by using InternVideo2 as the video encoder and train a video blip for communicating with open-sourced LLM Zheng et al. (2023); Jiang et al. (2023). In training, the video encoder will be updated. Detailed training recipts are in Li et al. (2023b).
@@
| Stage 1 | KMash | Web Video | 2M | Action Label |
@@
| Self-collected | Youtube Video | 60M | Generated Caption | |
@@ 4 Multimodal Video Data
We prepare our training data (video-only, video multimodal versions, and visual instructions) as given in Tab. 1.
@@ 4.1 Video-only Data for Masked AutoEncoders
We curate a video collection without labels and text named K-Mash from renowned action recognition datasets Carreira and Zisserman (2017); Goyal et al. (2017b); Monfort et al. (2020); Heilbron et al. (2015); Zhao et al. (2019), as detailed in Supp. It encompasses a wide range of video types, including first- and third-person perspectives, with both short and long durations, and featuring various characters and settings. Further, we give K-Mash with additionally sourced and meticulously selected 844K videos from YouTube for diversity. We observe notable action related task performance improvements in finetuned settings using K-Mash, as given in Sec. 5.4.2.
4.2 Videos with Audio-Speech Modalities for Contrastive Learning
We build a multimodal video dataset MVid with video-audio-speech information and their descriptions for pushing video perception via other modalities. It consists of 110M videos along with their VAS captions. In MVid, we collect videos from several sources (detailed in the supplementary material), segment them into clips, and automatically annotate them based on their solo unimodal or crossmodal inputs. We highlight the importance of temporal segmentation in clip generation and our video mutimodal annotation system. We find refining them leads to notable downstream improvements.
@@ 4.2 Videos with Audio-Speech Modalities for Contrastive Learning
We employ a temporal boundary detection model AutoShot Zhu et al. (2023c) to segment videos into clips instead of SceneDet filter from FFMPEG. It predicts boundaries based on temporal semantic variations rather than pixel differences, leading to generating semantically complete cuts without mixing extra frames with inconsistent context.
@@
We design a video multimodal annotation system VidCap to give proper unimodal and crossmodal descriptions for textualizing videos from different perceptions. It automatically captions visual, audio, and speech of MVid, then it corrects them and fuses them for cross-modal captions via LLM. The system frame is given in Fig. 3. VidCap has independent video, audio, speech captioner, and a LLM for caption refinement and fusion. For video, speech, and caption postprocessing, we employ existing methods as the video captioning pipeline in Wang et al. (2023d), WhisperV2-large model Radford et al. (2023), and Vicuna-1.5 Zheng et al. (2023). For audio, we craft a audio captioner upon VideoChat Li et al. (2023d), as we find no open-sourced one. It extracts audio features from inputs by Beats Chen et al. (2023d). We learn it by only tuning its Qformer using a combination of the large-scale audio-text corpus WavCaps Mei et al. (2023) dataset. Details are given in the supplementary material.
4.2.1 Video Instruction-Tuning Data for Next Token Prediction
We employ the training part of MVBench Li et al. (2023b). It comprises around 2M samples (both images and videos) from 34 distinct sources. This training dataset covers key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question anwser, 4) reasoning, and 5) classification.
@@ 5 Experiments
In our evaluation of InternVideo2, we assess the model performance of the three learning stages. The evaluation covers a wide range of tasks, including action recognition, video retrieval, question-answering, and more. It includes various scenarios such as zero-shot learning, finetuning, and others.
For InternVideo2 trained in stage 1, 2, and 3, we denote them with InternVideo2, InternVideo2, and InternVideo2, respectively. We also learn a CLIP-style InternVideo2 indicated by InternVideo2. It is post-pretrained from InternVideo2 by only preserving video and text encoders and contrastive loss.
Each training stage of InternVideo2-6B utilizes different configurations. In the first stage, we employ 256 NVIDIA A100 GPUs and train the model for 18 days. The second stage also utilizes 256 A100 GPUs and spans a training period of 14 days. Finally, in the third stage, we use 64 A100 GPUs and train the model for 3 days. We introduce DeepSpeed and FlashAttention Dao et al. (2022) for training and inference. Details and more results are given in the supp.
@@ 5.1.1 Action Recognition
We test InternVideo2 on Kinetics (i.e., K400, 600 and 700 Carreira and Zisserman (2017); Carreira et al. (2018, 2019)), Moments in Time V1 (MiT) Monfort et al. (2020), Something-Something V2 (SSv2) Goyal et al. (2017b), UCF Soomro et al. (2012), HMDB Kuehne et al. (2011), Charades Gao et al. (2017), ActivityNet Heilbron et al. (2015) (ANet) and HACS Zhao et al. (2019). We evaluate in four settings: (a) end-to-end finetuning the whole backbone; (b) attentive probing is similar to linear pooling, but extra trains the attention pooling layer Yu et al. (2022). (c) linear probing which freezes the backbone and only trains the task head; and (d) zero-shot.
| Method | Training Data | Setting | K400 | K600 | K700 | SthSthv2 | MiT | ANet | HACS |
@@
| MTV-H Wang et al. (2023a) | IV-370M | 32 280 | 89.9 | 90.3 | 83.4 | - | - | - | - |
| VideoMAEv2-g Wang et al. (2023a) | V-1.35M | 64 266 | 90.0 | 89.9 | - | 77.0 | - | - | - |
@@
| InternVideo2-1B | IV-1.1M | 8 224 | 91.3 | 91.4 | 85.0 | 77.1 | 50.8 | - | - |
| InternVideo2-1B | IV-1.1M | 16 224 | 91.6 | 91.6 | 85.4 | 77.1 | 50.9 | - | - |
| InternVideo2-6B | IV-2M | 8 224 | 91.9 | 91.7 | 85.7 | 77.5 | 51.0 | - | - |
| InternVideo2-6B | IV-2M | 16 224 | 92.1 | 91.9 | 85.9 | 77.4 | 51.2 | 95.9 | 97.0 |
@@
Tab. 2 shows InternVideo2-6B obtains new state-of-the-art (SOTA) results on Kinetics (92.1%/91.9%/85.9% on K400/600/700, respectively), SthSthv2, MiT, ANet, and HACS with only 16 frames, while the previous SOTAs require larger resolution (224 vs. 576) or model ensemble. As for MiT in Table 2, InternVideo2-6B exceeds the previous SOTA, CoCa-g, by a significant margin of 2.2% (51.2% vs. 49.0%). Regarding the temporal-related actions in Table 2, our InternVideo2-6B also surpasses MVD Wang et al. (2023b) on SSv2 (77.5% vs. 77.3%). Moreover, our InternVideo2-6B showcases top performance on untrimmed video analysis, as indicated in Table 2, with 95.9% on ActivityNet and 97.0% on HACS. These results affirm our model’s superior capability for robustly identifying complex actions across varied scenes.
@@
| VideoMAEv2-H Wang et al. (2023a) | 12 224 | 25.8 | - | 56.4 | 34.1 |
| TVTSv2-H Zeng et al. (2023b) | 12 224 | 73.1 | - | 91.8 | 65.7 |
@@
| InternVideo2-1B | 16 224 | 81.6 | 46.3 | 96.0 | 71.6 |
| InternVideo2-6B | 16 224 | 82.0 | 47.8 | 96.3 | 71.8 |
| InternVideo2-6B | 16 224 | 84.2 | 56.7 | 97.3 | 80.7 |
@@
| CLIP Radford et al. (2021a) | 12 | I-400M | 68.9 | 43.2 | - | 29.6 | - |
| TVTSv2 Zeng et al. (2023b) | 12 | V-8.5M | 78.0 | 52.1 | - | 48.4 | - |
@@
| InternVideo2-1B | 8 | IV-25.5M | 88.8 | 53.9 | 31.6 | 61.5 | 32.9 |
| InternVideo2-6B | 8 | IV-400M | 89.5 | 56.7 | 32.9 | 63.5 | 34.6 |
@@
| CLIP Radford et al. (2021a) | 8 | 58.4 | 70.1 | 55.1 | 67.2 | 46.1 | 58.4 |
| EVA-CLIP-L Sun et al. (2023c) | 1 | - | 65.0 | - | 64.9 | - | 59.1 |
| EVA-CLIP-E Sun et al. (2023c) | 1 | - | 69.8 | - | 69.3 | - | 63.4 |
@@
| InternVideo2-1B | 8 | 73.1 | 82.4 | 72.8 | 81.8 | 64.9 | 75.2 |
| InternVideo2-6B | 8 | 72.7 | 82.2 | 71.7 | 81.2 | 64.2 | 75.2 |
@@ 5.1.2 Temporal Action Localization
We evaluate models on four temporal action localization (TAL) datasets: THUMOS14 Idrees et al. (2017), ActivityNet Krishna et al. (2017), HACS Segment Zhao et al. (2019) and FineAction Liu et al. (2022) in a feature-based manner with finetuning. We employ output of the 7-th layer from InternVideo2 for inputs as the corresponding features without fusing anything else. ActionFormer Zhang et al. (2022) is used as the detection head. We report mean Average Precision(mAP) under multiple tIoU as in Lin et al. (2019); Chen et al. (2022b); Zhang et al. (2022); Yang et al. (2023). In Table 7, InternVideo2-6B gets the highest mAP among all comparisons in all datasets, while InternVideo2-1B nearly surpass other methods except in THUMOS14. We find InternVideo2-6B almost consistently improves mAP with a notable margin from InternVideo2-1B except in FineAction. We suppose scaling model capacity without data refinement cannot nontrivially improve fine-grained discrimination abilities of models. Scaling detailed annotations in training may address this issue.
@@ 5.1.3 Video Instance Segmentation
We evaluate on the popular Video Instance Segmentation (VIS) dataset, Youtube-VIS 2019 Yang et al. (2019). Built upon Mask2Former Cheng et al. (2021), we employ the video encoder of InternVideo2 as backbone with ViT-adapter Chen et al. (2022c) for features. We also try InternViT Chen et al. (2023a) for comparisons. In Table 8, InternVideo2 gets the highest mAP among all. It improves mAP with a nontrivial margin over InternViT and Swin-L. This validates its effectiveness in relatively fine-grained spatiotemporal perception.
5.2 Video-Language Tasks
@@ 5.2.1 Video Retrieval
We evaluate the video retrieval on six popular benchmarks Heilbron et al. (2015); Xu et al. (2016); Rohrbach et al. (2015); Anne Hendricks et al. (2017); Chen and Dolan (2011), as shown in Tab. 9 and 10. In eval, eight frames from the input videos are uniformly sampled. We report R@1 scores for both text-to-video (t2v) and video-to-text (v2t) tasks in Tab. 9 and 10. R@5 and R@10 are given in Supp.
@@
| Method | MSR-VTT | VATEX | MSVD | ANet Cap |
| BLIP2 Li et al. (2022b) | 19.0 | - | 27.4 | - |
| Video-LLaMA Zhang et al. (2023a) | 29.1 | - | 47.0 | - |
| Student Captioning Model Chen et al. (2024b) | 31.5 | - | 49.2 | - |
| VideoCoCa Yan et al. (2022) | 27.1 | 22.8 | - | 19.3 |
| Flamingo-3B Alayrac et al. (2022) | - | 40.1 | - | - |
| Flamingo-9B Alayrac et al. (2022) | - | 39.5 | - | - |
| VideoPrism-B Zhao et al. (2024) w/ PaLM-1B | 40.3 | 24.2 | - | - |
| VideoPrism-B Zhao et al. (2024) w/ PaLM-8B | 38.5 | 31.7 | - | - |
| InternVideo2-1B | 43.5 | 49.2 | 93.1 | 19.6 |
5.2.2 Video Captioning
As shown in Tab. 11,we evaluate zero-shot video captioning using InternVideo2 learned in stage 3 on MSR-VTT, VATEX, MSVD, and ANet Captions. InternVideo2 delivers superior CIDEr scores than VideoCoCa, Flamingo, and VideoPrism. Despite models scale, it may suggest a progressive learning scheme involving masked reconstruction, crossmodal contrastive learning, and next token prediction would present better video-to-text generation performance than those which only utilize some of them.
5.2.3 Multi-Choice Video Question Answering
We evaluate zero-shot multi-choice (MC) video QA using InternVideo2 in stage 2 on MSR-VTT, VATEX, and EgoSchema Mangalam et al. (2023). Note EgoSchema is a long-form egocentric video benchmark in the form of video QA. Tab. 12 shows InternVideo2 consistently improves MC accuracy compared with previous SOTAs except on LSMDC, where it gets a comparable result with InternVideo. Its zero-shot result on EgoSchema increases previous state-of-the-art by 9 points, suggesting its better generalization on long video understanding than others.

### v2

@@ 3.3 Stage3: Predicting Next Token with Video-Centric Inputs
To further enrich the semantics embedded in InternVideo2 and improve its support for video-centric dialogue, we tune it by connecting it to a LLM with QFormer design Li et al. (2022a, b). We employ the progressive learning scheme in Li et al. (2023d) by using InternVideo2 as the video encoder and train a video blip for communicating with open-sourced LLM Zheng et al. (2023); Jiang et al. (2023). Additionally, we implement a high-definition post-training stage to improve the model’s fine-grained and long spatiotemporal capabilities. During this stage, the input video is divided into up to six sub-videos with a resolution of 224x224 pixels each, along with one global resized sub-video of the same resolution. We then train the model for two additional epochs: the first epoch uses 8-frame video inputs, while the second epoch uses 16-frame inputs. During the additional training process, we update the video encoder and BLIP Qformer, while the LLM is updated using LoRA Hu et al. (2021).
@@
| Stage 1 | KMash | Web Video | 2M | - |
@@
| InternVid2 | Youtube Video | 50M | Generated Caption | |
@@ 4 Multimodal Video Data
We list our training data in Tab. 1. Among the datasets used, KMash and InternVid2 are newly built, whereas the rest are publicly available.
@@ 4.1 Video-only Data for Masked Autoencoders
We curate a new video set without labels named K-Mash from action recognition datasets Carreira and Zisserman (2017); Goyal et al. (2017a); Monfort et al. (2020); Heilbron et al. (2015); Zhao et al. (2019), as detailed in Supp. It encompasses a wide range of video types, including first- and third-person perspectives, with both short and long duration, and featuring various settings. Further, we give K-Mash2M with additionally sourced and selected 844K videos from YouTube for diversity.
4.2 Videos with Audio-Speech Modalities
We build a multimodal video dataset, coined as InternVid2, with video-audio-speech information and their descriptions for strengthening video perception via other modalities. It consists of 100M videos along with their VAS captions. In InternVid2, we collect videos from several sources (detailed in the supp), segment them into clips, and automatically annotate them based on their unimodal or crossmodal inputs. We highlight the importance of temporal segmentation in clip generation and our video mutimodal annotation system. We find refining them leads to notable downstream improvements.
@@ 4.2 Videos with Audio-Speech Modalities
We employ a temporal boundary detection model AutoShot Zhu et al. (2023c) to segment videos into clips instead of SceneDet filter from FFmpeg. It predicts boundaries based on temporal semantic variations rather than pixel differences, and is able to generate semantically complete cuts without mixing extra frames with inconsistent context.
@@
We design a video multimodal annotation system VidCap to give proper unimodal and crossmodal descriptions for textualizing videos from different perceptions. It automatically captions visual, audio, and speech of InternVid2, then it corrects them and fuses them for cross-modal captions via LLM. The system frame is given in Fig. 3. VidCap has independent video, audio, speech captioner, and a LLM for caption refinement and fusion. For video, speech, and caption postprocessing, we employ existing methods as the video captioning pipeline in Wang et al. (2023d), WhisperV2-large model Radford et al. (2023), and Vicuna-1.5 Zheng et al. (2023). For audio, we craft a audio captioner upon VideoChat Li et al. (2023c), as we find no open-sourced one. It extracts audio features from inputs by Beats Chen et al. (2023a). We learn it by only tuning its Qformer using a combination of the large-scale audio-text corpus WavCaps Mei et al. (2023) dataset. Details are given in the supplementary material.
4.3 Instruction-Tuning Data for Video Dialogue
We employ a updated training version of MVBench Li et al. (2023d). Originally, it comprises 1.9M samples (both images and videos) from 34 distinct sources. We decrease the amount of caption data from WebVid and CoCo to 80k and 100k separately and add new data from the S-MiT to boost the diversity rather than quantity of the instruction dataset. In the additional HD training stage, we incorporate the videos with corresponding GPT-4 annotations from Chen et al. (2024a) into the training set. We further expand the training set by including datasets from PerceptionTestQA Patraucean et al. (2024), TVQA Lei et al. (2018), NTU-RGB-D Liu et al. (2020), and EgotaskQA Grauman et al. (2022), along with grounding datasets based on DiDeMo Anne Hendricks et al. (2017b) and COCO Lin et al. (2014). This training data encompasses key features of image and video understanding across crucial tasks, including 1) conversation, 2) caption, 3) visual question answer, 4) reasoning, and 5) classification.
@@ 5 Experiments
In our evaluation of InternVideo2, we assess the models from three learning stages. The evaluation covers a wide range of tasks, including video recognition, video retrieval, question-answering, and more. It includes various scenarios such as zero-shot learning, finetuning, and linear probing.
For InternVideo2 trained in stage 1, 2, and 3, we denote them with InternVideo2s1, InternVideo2s2, and InternVideo2s3, respectively. We also learn a CLIP-style InternVideo2 indicated by InternVideo2clip. It is post-pretrained from InternVideo2s2 by only preserving video and text encoders and contrastive loss.
Each training stage of InternVideo2-6B uses different configurations and resources. In the first stage, we employ 256 NVIDIA A100 GPUs and train the model for 18 days. The second stage also utilizes 256 A100 GPUs and spans a training period of 14 days. Finally, in the third stage, we use 64 A100 GPUs and train the model for 3 days. We introduce DeepSpeed and FlashAttention Dao et al. (2022) for training and inference. More implementation details and experiment results are given in the supp.
@@ 5.1.1 Action Recognition
We test InternVideo2 on Kinetics (i.e., K400, 600 and 700 Carreira and Zisserman (2017); Carreira et al. (2018, 2019)), Moments in Time V1 (MiT) Monfort et al. (2020), Something-Something V2 (SSv2) Goyal et al. (2017a), UCF Soomro et al. (2012), HMDB Kuehne et al. (2011), Charades Gao et al. (2017), ActivityNet Heilbron et al. (2015) (ANet) and HACS Zhao et al. (2019). We evaluate in four settings: (a) end-to-end finetuning the whole backbone; (b) attentive probing is similar to linear pooling, but extra trains the attention pooling layer Yu et al. (2022). (c) linear probing which freezes the backbone and only trains the task head; and (d) zero-shot.
| Method | Training Data | Setting | K400 | K600 | K700 | Sth-Sthv2 | MiT | ANet | HACS |
@@
| MTV-H Wang et al. (2023b) | IV-370M | 32 280 | 89.9 | 90.3 | 83.4 | - | - | - | - |
| VideoMAEv2-g Wang et al. (2023b) | V-1.35M | 64 266 | 90.0 | 89.9 | - | 77.0† | - | - | - |
@@
| InternVideo2s1-1B | IV-1.1M | 8 224 | 91.3 | 91.4 | 85.0 | 77.1 | 50.8 | - | - |
| InternVideo2s1-1B | IV-1.1M | 16 224 | 91.6 | 91.6 | 85.4 | 77.1 | 50.9 | - | - |
| InternVideo2s1-6B | IV-2M | 8 224 | 91.9 | 91.7 | 85.7 | 77.5 | 51.0 | - | - |
| InternVideo2s1-6B | IV-2M | 16 224 | 92.1 | 91.9 | 85.9 | 77.4 | 51.2 | 95.9 | 97.0 |
@@
Tab. 2 shows InternVideo2-6B obtains new state-of-the-art (SOTA) results on Kinetics (92.1%/91.9%/85.9% on K400/600/700, respectively), SthSthv2, MiT, ANet, and HACS with only 16 frames, while the previous SOTAs require larger resolution (224 vs. 576) or model ensemble. As for MiT in Table 2, InternVideo2-6B exceeds the previous SOTA, CoCa-g, by a significant margin of 2.2% (51.2% vs. 49.0%). Regarding the temporal-related actions in Table 2, our InternVideo2-6B also surpasses MVD Wang et al. (2023c) on SSv2 (77.5% vs. 77.3%). Moreover, our InternVideo2-6B showcases top performance on untrimmed video analysis, as indicated in Table 2, with 95.9% on ActivityNet and 97.0% on HACS. These results affirm our model’s superior capability for robustly identifying complex actions across varied scenes. Note “I" and “V" denotes images and videos, respectively. “IV-3B" means the total number of the used images and videos is 3B, while “I-3B" means using 3B images.
@@
| VideoMAEv2-H Wang et al. (2023b) | 12 224 | 25.8 | - | 56.4 | 34.1 |
| TVTSv2-H Zeng et al. (2023) | 12 224 | 73.1 | - | 91.8 | 65.7 |
@@
| InternVideo2s1-1B | 16 224 | 81.6 | 46.3 | 96.0 | 71.6 |
| InternVideo2s1-6B | 16 224 | 82.0 | 47.8 | 96.3 | 71.8 |
| InternVideo2s2-6B | 16 224 | 84.2 | 56.7 | 97.3 | 80.7 |
@@
| CLIP Radford et al. (2021b) | 12 | I-400M | 68.9 | 43.2 | - | 29.6 | - |
| TVTSv2 Zeng et al. (2023) | 12 | V-8.5M | 78.0 | 52.1 | - | 48.4 | - |
@@
| InternVideo2clip-1B | 8 | IV-25.5M | 88.8 | 53.9 | 31.6 | 61.5 | 32.9 |
| InternVideo2clip-6B | 8 | IV-400M | 89.5 | 56.7 | 32.9 | 63.5 | 34.6 |
@@
| CLIP Radford et al. (2021b) | 8 | 58.4 | 70.1 | 55.1 | 67.2 | 46.1 | 58.4 |
| EVA-CLIP-L Sun et al. (2023b) | 1 | - | 65.0 | - | 64.9 | - | 59.1 |
| EVA-CLIP-E Sun et al. (2023b) | 1 | - | 69.8 | - | 69.3 | - | 63.4 |
@@
| InternVideo2clip-1B | 8 | 73.1 | 82.4 | 72.8 | 81.8 | 64.9 | 75.2 |
| InternVideo2clip-6B | 8 | 72.7 | 82.2 | 71.7 | 81.2 | 64.2 | 75.2 |
@@ 5.1.2 Temporal Action Localization
We evaluate models on four temporal action localization (TAL) datasets: THUMOS14 Idrees et al. (2017), ActivityNet Krishna et al. (2017), HACS Segment Zhao et al. (2019) and FineAction Liu et al. (2022) in a feature-based manner with finetuning. We employ output of the 7-th layer from InternVideo2 for inputs as the corresponding features without fusing anything else. ActionFormer Zhang et al. (2022) is used as the detection head. We report mean Average Precision(mAP) under multiple tIoU as in Lin et al. (2019); Chen et al. (2022a); Zhang et al. (2022); Yang et al. (2023). In Table 7, InternVideo2-6B gets the highest mAP among all comparisons in all datasets, while InternVideo2-1B nearly surpass other methods except in THUMOS14. We find InternVideo2-6B almost consistently improves mAP with a notable margin from InternVideo2-1B except in FineAction. We suppose scaling model capacity without data refinement cannot nontrivially improve fine-grained discrimination abilities of models. Scaling detailed annotations in training may address this issue.
@@ 5.1.3 Video Instance Segmentation
We evaluate on the Video Instance Segmentation (VIS) dataset Youtube-VIS 2019 Yang et al. (2019). Built upon Mask2Former Cheng et al. (2021), we employ the video encoder of InternVideo2 as backbone with ViT-adapter Chen et al. (2022b) for features. We also try InternViT Chen et al. (2023c) for comparisons. In Table 8, InternVideo2 gets the highest mAP among all. This validates its effectiveness in relatively fine-grained spatiotemporal perception.
5.2 Video-Audio-Language Tasks
@@ 5.2.1 Video Retrieval
We evaluate the video retrieval on six popular benchmarks Heilbron et al. (2015); Xu et al. (2016); Rohrbach et al. (2015); Anne Hendricks et al. (2017a); Chen and Dolan (2011), as shown in Tab. 9 and 10. In eval, eight frames from the input videos are uniformly sampled. We report R@1 scores for both text-to-video (t2v) and video-to-text (v2t) tasks in Tab. 9 and 10. R@5 and R@10 are given in Supp.